OpenStudy (anikate):

Decide whether the sequence is an arithmetic, Geometric, or neither. Explain your answer. 6,24,96,384

OpenStudy (anikate):

@Calcmathlete

OpenStudy (anikate):

what does arithmetic and geometric mean?

OpenStudy (anonymous):

i think the answer would be arithmetic?

OpenStudy (anonymous):

Well, arithmetic sequences have a common difference. So if \(a_2 - a_2\) remains constant when doing \(a_3 - a_2\) and so on, it's arithmetic. Geometric sequences have a common ratio. So if \(a_2/a_1\) remains constant when doing \(a_3/a_2\) and so on, then it's geometric. If it's neither of these, then it's neither.

OpenStudy (anonymous):

So what do you think it is?

OpenStudy (anikate):

i dont understand ur definition.

OpenStudy (anikate):

what is a2-a1

OpenStudy (anonymous):

\(a_2\) is the second term of the sequence.

OpenStudy (anonymous):

Let's see if looking at an example will help. 1, 2, 4, 8, 16... This is a geometric sequence. 1, 2, 3, 4, 5, 6, 7... THis is an arithmetic sequence.

OpenStudy (anikate):

and what do you mean by "if a2−a2 remains constant when doing a3−a2"

OpenStudy (anonymous):

So let's take a look. \[a_2 - a_1 \implies 24 - 6 \implies 18\]Now let's do \(a_3 - a_2\) \[a_3 - a_2 \implies 96 - 24 \implies 72\]\[18 ≠ 72\]

OpenStudy (anonymous):

Therefore, it's not an arithmetic sequence.

OpenStudy (anonymous):

Now let's test to see if it's a geometric sequence. \[a_2/a_1 \implies 24/6 \implies 4\]\[a_3/a_2 \implies 96/24 \implies 4\]\[4 = 4\]Therefore, it's a geometric sequence. Get it?

OpenStudy (anikate):

where did u get 24 from?

OpenStudy (anonymous):

24 is the second term in the sequence.

OpenStudy (anikate):

oh ok rite MY SEQUENCE not ur example

OpenStudy (anonymous):

lol yeah.

OpenStudy (anikate):

wait why did u do a2-a1. is that the formula?

OpenStudy (anonymous):

THat's just how you can test for the common difference. You can honestly do that for any of the terms as long it's a term - the term previous.

OpenStudy (anonymous):

Honestly though, you can see just by looking at a sequence if it's geometric or arithmetic. For instance. If I gave you the choice to say arithmetic or geometric for the following sequence, which one would you think? 1, 5, 25, 125, 62...

OpenStudy (anikate):

sry can u plz explain those same steps again my copy and pasting

OpenStudy (anikate):

@Topkart33 com here

OpenStudy (anonymous):

Alright. To find a common difference which can also test to see if the sequence is arithmetic, just do \(a_2 - a_1\) or \(a_3 - a_2\) or \(a_4 - a_3\) and so on... To find a common ratio which can also test to see if the sequence is geometric, just do \(a_2/a_1\) or \(a_3/a_2\) and so on...Once you get used to working with series and sequences, you'll do this instinctively.

OpenStudy (anikate):

what's the common difference?

OpenStudy (anonymous):

You mean a definition?

OpenStudy (anikate):

yeah

OpenStudy (anikate):

calcmath u still there? plz dont give me a link unless it has pictures to explain

OpenStudy (anonymous):

Alright. A common difference is a number in arithmetic sequences that when added to terms of the sequence gives you the following term.

OpenStudy (anikate):

so it's only in arithmetic sequences? I dont understand that definition

OpenStudy (anikate):

and what terms are you talking about in ur definition?

OpenStudy (anonymous):

It's only in arithmetic sequences. Terms are the numbers in a sequnce or series. In your sequence, the terms are 6, 24, 96, 384

OpenStudy (anikate):

what do u mean by following term in your definition?

OpenStudy (anikate):

Alright. A common difference is a number in arithmetic sequences that when added to terms of the sequence gives you the following term. /i\ i i what is the following term?

OpenStudy (anonymous):

The term that comes after what you added the term to. If I add the common difference to \(a_1\), then you should get \(a_2\).

OpenStudy (anikate):

ok what's the common difference in my sequence: 6,24,96,384

OpenStudy (anikate):

calcmath...hello?

OpenStudy (anonymous):

I am here. There is no common difference there since the sequence is geometric...

OpenStudy (anikate):

ok now next question: -11, -7,-3,1,.....

OpenStudy (anikate):

I will answer it

OpenStudy (anikate):

u tell me if its right of wrong only

OpenStudy (anikate):

is it arithmetic?

OpenStudy (anonymous):

Yup. What do you think the common difference is?

OpenStudy (anikate):

idk? can u show me?

OpenStudy (anikate):

calc

OpenStudy (anonymous):

Do you at least have a guess?

OpenStudy (anikate):

no I dont know how to figure it out. my guess is -4

OpenStudy (anonymous):

It would be a positive 4. \[-7 - (-11) = 4\]\[-3 - (-7) = 4\]\[1 - (-3) = 4\]See? They all say that it's 4 :)

OpenStudy (anikate):

ok thx can you plz stay on this chat cause I am going to answer more questions and I need you to check it for me then I need help on other questions.

OpenStudy (anikate):

-3/5, 4/25, 5/125, 6/625

OpenStudy (anonymous):

Alright.

OpenStudy (anikate):

dont anwer it

OpenStudy (anikate):

I will answer it, you will check my answer :) ?

OpenStudy (anikate):

not to be rude...

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

Well...

OpenStudy (anikate):

is it neither?

OpenStudy (anikate):

It took alwhile becasue I was working it out

OpenStudy (anonymous):

Yup. It's neither.

OpenStudy (anikate):

when I am not typing, dont think that I am not on :D I am always on. I think it's rude to leave someone waiting for you

OpenStudy (anonymous):

Yeah...if I don't respond in under a minute, I'm still here, so don't freak out lol

OpenStudy (anikate):

alright next question: 4, 13, 22, 31

OpenStudy (anikate):

is it arithmetic?

OpenStudy (anonymous):

Yes. And the common difference?

OpenStudy (anikate):

9

OpenStudy (anonymous):

Yes!

OpenStudy (anikate):

YAY! next question: 1/3, 2/3, 1, 4/3

OpenStudy (anikate):

is it arichmetic?

OpenStudy (anonymous):

Yes.

OpenStudy (anikate):

alright and common difference is 1/3

OpenStudy (anikate):

umm calc?

OpenStudy (anikate):

calc?

OpenStudy (anonymous):

Yes. You are correct.

OpenStudy (anikate):

alright next

OpenStudy (anikate):

Find the common ratio of the geometric sequence. 3, 6, 12, 24

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

Do you know how to do it?

OpenStudy (anikate):

i dont know how to do it?

OpenStudy (anonymous):

Alright. \[6/3 = 2\]\[12/6 = 2\]\[24/12 = 2\]

OpenStudy (anikate):

umm so the common ratio is 2?

OpenStudy (anonymous):

Yes.

OpenStudy (anikate):

oooh

OpenStudy (anikate):

ok next question. 5, 40, 320, 2560, ....

OpenStudy (anonymous):

Can you try this one?

OpenStudy (anikate):

YES WAIT i AM DOING IT :D

OpenStudy (anikate):

8?

OpenStudy (anonymous):

Yup :)

OpenStudy (anikate):

alright next: -1/4, 1/8, -1/16, 1/32

OpenStudy (anikate):

-1/2

OpenStudy (anonymous):

Yup :)

OpenStudy (anikate):

alright thx next: Find the 10th term and write a rule for the nth term of the geometric sequence then find asub6: 6, -30, 150, -750

OpenStudy (anonymous):

Are you familiar with these?

OpenStudy (anikate):

nope how do I do them?

OpenStudy (anonymous):

Ok. For these, first find what the common ratio is.

OpenStudy (anikate):

-5

OpenStudy (anonymous):

Ok. Now are you familiar with the formulas \[a_n = a_1 + (n - 1)d\]and\[a_n = a_1r^{n - 1}\]

OpenStudy (anikate):

I remember those now but need help using them

OpenStudy (anikate):

what is the d, n, and r

OpenStudy (anonymous):

d = common difference r = common ratio n = number of terms